OpenStudy (anonymous):

A cubic box is 5 inches on each side. If each dimension is increased by x inches, What is the polynomial function that models the new volume V? (V=LWH)

OpenStudy (sw050399):

It's a cube, so you know all sides are equal. If x inches is added to all sides, you have x + 5 for each side. This means that L = x + 5, H = x + 5, and W = x + 5. So, if you want the volume plug the values in the formula you have above and that will be your answer.

OpenStudy (anonymous):

Thanks a lot!!